A quadratic equation f (x) has two roots α and β . If α=-4+5i, determine the root β
and thus obtain an expression for the equation f (x)
1
Expert's answer
2020-06-08T19:29:55-0400

If "\alpha=-4+5i" is the root of the quadratic equation "f(x)=0," then "\beta=-4-5i" is also the root of the quadratic equation "f(x)=0."

"\alpha=-4+5i, \beta=-4-5i"

Hence


"f(x)=a(x-\alpha)(x-\beta), a\not=0"

"f(x)=a(x-(-4+5i))(x-(-4-5i))"

"\alpha+\beta=-4+5i-4-5i=-8=-{b\over a}"

"\alpha\cdot\beta=(-4+5i)\cdot(-4-5i)=(-4)^2+(5)^2=41={c\over a}"

"f(x)=ax^2 +8ax+41a, a\not=0"